Find the explicit formula for the geometric sequence with a9 = 128 and r = 2.
Answer by Cromlix(4381)

Hi there,
Using the formula:
a(n) = a(1)(r)^n -1
a(9) = a(1)*(2)^8 = 128
a(1)* 256 = 128
a(1) = 0.5
Explicit formula =
a(n) =0.5*(r)^n-1
Hope this helps:-)